Consider a bond (with par value = $1,000) paying a coupon rate of 7% per year semiannually when the market interest rate is only 6% per half-year. The bond has 3 years until maturity.

a. Find the bond's price today and 6 months from now after the next coupon is paid.

b. What is the total (6-month) rate of return on the bond?

Hint: rate of return = (Interest amount + Price Appreciation)/Initial Price
Answer #1

a. Find the bond's price today
=(1000*7%/2)/(6%)*(1-1/1.06^6)+1000/1.06^6=877.0668918

price 6 months from now after the next coupon is paid.
=(1000*7%/2)/(6%)*(1-1/1.06^5)+1000/1.06^5=894.6909054

b. What is the total (6-month) rate of return on the bond?
=(7%*1000/2+894.6909054)/877.0668918-1=6.00%